OpenStudy (anonymous):

The distance of a ball thrown upward is given by h(t) = -2t^2 + 12t – 10 meters. Find the velocity of the ball when it returns to the ground.

OpenStudy (anonymous):

x=2t^2+12t-10 dx/dt=4t+12 t=0 in above equation.. 12ms^-1

OpenStudy (anonymous):

differentiate

OpenStudy (anonymous):

i dont know if my answer is correct...sorry if its wrong

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

I think it is. I got that before and its definitely wrong sweetie :/

OpenStudy (anonymous):

hey :) can anyone help me out? :/

OpenStudy (anonymous):

-4t+12 =12ms^-1

OpenStudy (anonymous):

@rhino..assuming that the ball is thrown at time t=0..we would get velocity of the ball when it hits the earth wouldnt we?shouldnt we consider the acceleration..??sorry if its nonsense

OpenStudy (anonymous):

when on the ground t is 0 -4(0)+12 =12

OpenStudy (anonymous):

are you sure it might not be 3,5,--8, or -10. I got 12 too but it seems to be wrong :( ughh

OpenStudy (anonymous):

velocity cant be negative..becausse its a vector...

OpenStudy (anonymous):

Thanks for trying everyone. :(

OpenStudy (anonymous):

veloicty CAN be negative because it is vector

OpenStudy (anonymous):

-4t+12=0 -4t=-12 t=3 -4(6)+12 -24+12=-12 m/s

OpenStudy (anonymous):

hey.... magnitude of vector is always positive....

OpenStudy (anonymous):

\[v(t)=-4t+12\]The ball comes to a rest at it's maximum height. Entering v=0:\[-4t=-12, t=3\]How high is the ball? We enter t=3 into position function:\[s(t)=-2(3)^2+12(3)-10=-18+36-10=8\]The ball will take the same amount of time to reach the ground as it did to reach the max height. The total time to go up and then back down is 6 sec. Inputting:\[v(6)=-4(6)+12=-12m/s\] This makes sense. The ball was released at an initial velocity of +12m/s (input t=0 into velocity function), comes to a stop under the action of gravity and reverses direction. When it reaches the same initial point it must have the same velocity as the initial velocity only in the opposite direction.

OpenStudy (anonymous):

you are correct, magnitude is always positive but we are not trying to find magnitude

OpenStudy (anonymous):

@imran ..The magnitude of a vector is always treated as non negative and the minus sign indicates the reversal of that vector through an angle of 180 degree. http://wiki.answers.com/Q/Can_the_magnitude_of_a_vector_have_a_negative_value

OpenStudy (anonymous):

magnitude of velocity is speed; the question ask for velocity;not speed

OpenStudy (anonymous):

read my previous post

OpenStudy (anonymous):

we are not finding magnitude, thus we don't need to worry about that

OpenStudy (anonymous):

then what r we finding?

OpenStudy (anonymous):

velocity which is a vector; not magnitude

OpenStudy (anonymous):

magnitude is scalar

OpenStudy (anonymous):

hey....vector has mag and dir... when u say he force is 20N due east ..u mean both dir and mag...even if u say the force is 20N it doesnt make force scalar..

OpenStudy (anonymous):

exactly vector has both magnitude and direction, so in our case 12 is magnitude - is direction; which makes vector -12

OpenStudy (anonymous):

moreover...u know representaion of vector?its represented using directed line segment and the length of the line segment gives magnitude...so length cant be negative...hence magnitude of vector cannot also be negative

OpenStudy (anonymous):

I agree with you that magnitude can't be negative, which is absolutely correct. What I am saying is vector can be negative; not its magnitude

OpenStudy (anonymous):

the answer is -8 :) thanks for trying everyone. -4(5) +12

OpenStudy (anonymous):

Velocity can in fact be negative. Think outside the box d: if im going 30 miles an hour east but you're going 80 miles an hour west you're going to be going faster but in a different direction

OpenStudy (anonymous):

Ah, I know why. The ball starts out at a position of x=-10m. If you define the "ground" to be x=0m than the answer is -8m/s as you say. I gave you the velocity when the ball returns to it's initial position of -10m NOT 0m.

OpenStudy (anonymous):

oh yea the ball is coming down...... and hey why 5?

OpenStudy (anonymous):

Hit the ground at t=5

OpenStudy (anonymous):

Exactly :D

OpenStudy (anonymous):

The question should read "find the velocity of the ball when it returns to h=0m"

OpenStudy (anonymous):

I took the question as "find the velocity of the ball when it returns to it's initial position h=-10m"